OpenStudy (anonymous):

Simplify the radical √a^3

OpenStudy (anonymous):

a^3/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

a^(1.5)=a*sqrta

OpenStudy (anonymous):

a√a a a√a^2 a^2

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

first option is the answer

OpenStudy (austinl):

\(\sqrt{a^3}=\sqrt{a^2}\times\sqrt{a}=a\times\sqrt{a}=a\sqrt{a}\)

OpenStudy (debbieg):

Just break down what is under the radical into a product of the part that is a perfect square (anything that can be written with an even exponent), and the part that isn't. so \(\Large a^3=a^2 \cdot a\) and the square root of the perfect square "piece" gets to come outside the radical, the stuff that isn't a perfect square has to stay where it is. :)